OpenStudy (anonymous):

Geometry Reason and Proof can please check my answers? 1.One way to show that a statement is not a good definition is to find a ______. A.converse B. conditional c.biconditional d.counterexample*** my answer 3. Is the following definition of supplementary reversible? If yes write it as a true biconditional. Linears pairs of angles are supplementary. A. The statement is not reversible. B.Yes, if angles form a linear pair, then they are supplementary. C.Yes, angles form a linear pair if ( and only if) they are supplementary*** my answer D. Yes, angles are supplementary if they form a lin

OpenStudy (anonymous):

linear pair for the end of D it got marked off.

OpenStudy (anonymous):

looks good

OpenStudy (anonymous):

Thanks mind taking a look at one more? It's a bit more tricky.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Name the Property of Congruence that justifies the statement: If m<A+m<B=m<C,then m<A=m<C -m<B choices are (Transitive Property)(Symmetric Property and Reflexive Property or none of the above I think it's transitive.

OpenStudy (anonymous):

a + b = c a = c - b They subtracted b from both sides. that is the Subtraction Property of Equality Transitive would be "If A = B and B = C, then A = C." It's transferring equality between different terms.

OpenStudy (anonymous):

So it would be none of the above? I'm seeing...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright thanks,

OpenStudy (anonymous):

yw